OpenStudy (anonymous):

Write the slope-intercept form of the equation. 3x – 2y = -12

OpenStudy (anonymous):

slop =-3/2

OpenStudy (anonymous):

intercept=-6

OpenStudy (amistre64):

simply "solve for y" to get it into y = ________

OpenStudy (anonymous):

slope is given by y=bx+c

OpenStudy (anonymous):

where bx is your slope and c is your intercept

OpenStudy (anonymous):

would slope intercept form be -3/2?@emekaojei

OpenStudy (anonymous):

don"t get your question

OpenStudy (anonymous):

what would the slope intercept form of 3x – 2y = -12 be

OpenStudy (anonymous):

@emekaojei

OpenStudy (anonymous):

slop is a different thing and intercept too. unless you want to multiply the slop with intercept, then your answer will be 9

OpenStudy (anonymous):

are you clear now?

OpenStudy (anonymous):

just take your slope to be 3/2, your intercept to be 6 and slope intercept to be 9

OpenStudy (anonymous):

are you clear now?

OpenStudy (anonymous):

any option

OpenStudy (anonymous):

y intercept is 6, x intercept is -4, the slop is 3/2